OpenStudy (anonymous):

Solve the system of linear equations below: x - 3y = -3 x + 3y = 9

OpenStudy (anonymous):

A) x = -12, y = 7 B) x = 3, y = 2 C) x = 6, y = 1 D) x = 6, y = 2

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Can you explain it please

OpenStudy (igreen):

@otakukid99 Don't give answers.

OpenStudy (igreen):

We can solve by substitution. First, we can rearrange the first equation. \(x - 3y = -3\) Now we add 3y to both sides: \(x = -3 + 3y\) Okay, now we can plug in \(-3 + 3y\) for \(x\) in the 2nd equation. \(x + 3y = 9\) \(-3 + 3y + 3y = 9\) Can you simplify that? @TheYoungBlaze

OpenStudy (igreen):

We're solving for \(y\).

OpenStudy (anonymous):

-3 + 6y = 9 Add 3 to both sides to get rid of the -3 So 6y = 12 Divide by 6 1y = 2 Right? Or no?

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

So \(y = 2\). Now we can plug in 2 for y into any of the two original equations: \(x + 3y = 9\) \(x + 3(2) = 9\) Can you solve for \(x\)?

OpenStudy (igreen):

@TheYoungBlaze

OpenStudy (anonymous):

x + 6 = 9 6x = 9 Divide by GCF which is 3 2x = 3

OpenStudy (igreen):

Um..no. \(x + 6 = 9\) Just subtract 6 to both sides..what do you get? @TheYoungBlaze

OpenStudy (anonymous):

Ah I have a tendency to overthink things. I knew the answer sounded wrong. x = 3 So: y = 2 x = 3

OpenStudy (igreen):

Yep.

OpenStudy (igreen):

Just remember, we do the opposite..\(x + 6 = 9\). Notice that 6 is being ADDED to \(x\). To undo addition, we do the opposite of addition, which is subtraction..not Division or Multiplication.